can you please help me to figure out how to solve this
Caribbean paradise has 35 more suits than pacific paradise. med paradise has 37 fewer than twice the suits of pacific paradise. find the number of suits for each ship total is 664

Answer by JulietG(1812)

Let C be Caribbean, M be Mediterranean, and P be Pacific.
C + M + P = 664
C = P + 35
M = 2P - 37
Substitute the values of C and M from the bottom equations into the top.
C + M + P = 664
(P+35) + (2P-37) + P = 664
Add
4P - 2 = 664
Add 2 to each side
4P = 666
Divide each side by 4
P = 166.5
If P = 166.5, then C = 201.5, and M = 296

Check the numbers (or maybe my math is off??). It's odd that the ships have "half" a suit.